# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v1

@@ Abstract
Parameter-efficient fine-tuning for continual learning (PEFT-CL) has shown promise in adapting pre-trained models to sequential tasks while mitigating catastrophic forgetting problem. However, understanding the mechanisms that dictate continual performance in this paradigm remains elusive. To tackle this complexity, we undertake a rigorous analysis of PEFT-CL dynamics to derive relevant metrics for continual scenarios using Neural Tangent Kernel (NTK) theory. With the aid of NTK as a mathematical analysis tool, we recast the challenge of test-time forgetting into the quantifiable generalization gaps during training, identifying three key factors that influence these gaps and the performance of PEFT-CL: training sample size, task-level feature orthogonality, and regularization. To address these challenges, we introduce NTK-CL, a novel framework that eliminates task-specific parameter storage while adaptively generating task-relevant features. Aligning with theoretical guidance, NTK-CL triples the feature representation of each sample, theoretically and empirically reducing the magnitude of both task-interplay and task-specific generalization gaps. Grounded in NTK analysis, our approach imposes an adaptive exponential moving average mechanism and constraints on task-level feature orthogonality, maintaining intra-task NTK forms while attenuating inter-task NTK forms. Ultimately, by fine-tuning optimizable parameters with appropriate regularization, NTK-CL achieves state-of-the-art performance on established PEFT-CL benchmarks. This work provides a theoretical foundation for understanding and improving PEFT-CL models, offering insights into the interplay between feature representation, task orthogonality, and generalization, contributing to the development of more efficient continual learning systems.
@@ 1 Introduction
In practical applications, the relentless evolution of environments underscores the urgency for learning systems that can progressively accumulate knowledge. This has led to the prominence of Continual Learning (CL) [54, 87, 13, 60, 43, 78, 53], a cornerstone task that equips the learning models with the ability to seamlessly assimilate fresh information over time, while mitigating catastrophic forgetting, i.e., a phenomenon that erodes previously acquired knowledge. In recent years, with the proliferation of pre-trained models possessing strong generalization capabilities [6, 63], researchers have discovered that they can empower early exploratory methods [7, 80, 18, 22, 52, 4, 68, 97, 69, 90, 89, 99, 44, 46, 45, 33, 88], enabling CL systems to integrate new knowledge more efficiently. However, full fine-tuning of pre-trained models is computationally intensive and may compromise their original generalization capabilities [27, 50, 94]. Thus, as a promising paradigm, Parameter-Efficient Fine-Tuning for Continual Learning (PEFT-CL) emerges as an alternative, updating only a minimal set of additional parameters while keeping the pre-trained model intact. Specifically, PEFT-CL not only offers a more philosophically sound framework akin to Socratic dialogue [95] but also provides a lightweight training process that avoids generalization deterioration associated with full-scale fine-tuning [35, 76]. In addition, this seamless integration of new and old knowledge aligns with the wisdom expressed by Bernard of Chartres, demonstrating how PEFT-CL builds upon pre-existing knowledge to achieve a more adaptive learner with robust memory capabilities.
@@
Despite initial successes in mitigating catastrophic forgetting [82, 81, 71, 19, 100], PEFT-CL largely relies on subjective human insights and experiential doctrines for network design and enhancement, lacking a rigorous mathematical foundation. This reliance on non-theoretical approaches constrains the potential for a deeper understanding and advancement of the fundamental mechanisms within these learning systems. While Hide-Prompt [77] acknowledges the importance of addressing this issue and offers a loss-based perspective, it falls short of modeling optimization dynamics and pinpointing key factors. Therefore, to address this gap, we adopt the Neural Tangent Kernel (NTK) theory [32, 8, 5] as a robust mathematical tool to delve deeply into the intricacies of PEFT-CL optimization. Through this rigorous analysis, we derive several fundamental theorems and lemmas, including theorem 1, theorem 2, lemma 3, and theorem 4. While initially considered from a CL perspective, these have been generalized to the PEFT-CL scenario, providing profound insights into the key factors essential for effectively combating catastrophic forgetting in PEFT-CL optimization. Guided by these theories and key factors, we develop an NTK-CL framework, effectively reducing the quantified catastrophic forgetting discussed later.
@@
In addition to theoretical advantages, we also detail the differences in structure and optimization between our NTK-CL framework and current mainstream methodologies in Fig. 1. Unlike the Additional Subnetworks paradigm (Fig. 1a), which constructs task-specific subnetwork parameter spaces and concatenates features from all network parameter spaces at inference time [100, 19, 48], or the Prompts Optimization paradigm (Fig. 1b), which builds task-specific prompt pools for input interaction and employs cosine similarity for prompt selection [82, 36, 81, 71, 62], NTK-CL (Fig. 1c) eliminates the need for task-specific parameter storage or prompt pools. Instead, it leverages a shared network parameter space across all tasks to adaptively generate task-relevant features based on input characteristics. Specifically, its design and optimization are entirely derived from NTK-based generalization gaps, which not only triple the expansion of sample representations but also consider knowledge retention, task-feature dissimilarity, and regularization term.
@@
  • Theoretical Exploration of PEFT-CL: We pioneer the analysis of PEFT-CL through NTK lens and foundational mathematics. Through a series of derived theorems and lemmas, we identify critical factors that optimize PEFT-CL learners, including the number of samples in data subsets, the total sample volume across the dataset, knowledge retention strategies, task-feature dissimilarity constraints, and adjustments to regularization terms.
  • Innovative Solutions Based on Key Factors: Leveraging theoretical insights, we propose an NTK-CL framework, incorporating multiple interventions to expand each sample’s representational breadth, an adaptive Exponential Moving Average (EMA) mechanism that preserves intra-task NTK forms (enhancing knowledge retention), task-feature orthogonality constraints that attenuate inter-task NTK forms (increasing knowledge separability), and tailored regularization adjustments that meet the theoretical prerequisites for solving PEFT-CL dynamics. With these strategies, we optimally minimize the generalization gaps and population losses in both task-interplay and task-specific settings for PEFT-CL scenario, significantly mitigating the catastrophic forgetting problem both theoretically and practically.
  • Empirical Validation on Diverse Datasets: We conduct extensive experiments across various datasets to validate the effectiveness of our key factors and methodologies. Additionally, we perform fair comparisons against numerous state-of-the-art methods, ensuring consistent task segmentations to mitigate performance discrepancies. This comprehensive validation substantiates the efficacy of our theoretical innovations in practical applications.
@@ 2 Related Works
Parameter-Efficient Fine-Tuning has emerged as a pivotal paradigm for optimizing model performance while mitigating computational and memory burdens associated with large-scale model adaptation. Seminal works introduce diverse methodologies, including Adapter modules [26], Low-Rank Adaptation (LoRA) [27], Prefix Tuning [47], Prompt Tuning [6], and BitFit [96]. These approaches demonstrate the efficacy of selectively fine-tuning components or introducing compact, trainable sub-networks within pre-trained architectures. Subsequent advancements further expand PEFT’s scope and capabilities. Jia et al. [34] pioneer efficient prompt tuning techniques for vision transformers, extending PEFT’s applicability to the visual domain. Zhou et al. [102] introduce contextual prompt fine-tuning, enhancing model adaptability while preserving generalization. Recent comprehensive studies [20, 85, 86, 15] reinforce PEFT’s critical role in enhancing model generalization and efficiency. These investigations rigorously analyze the theoretical underpinnings and empirical efficacy of various PEFT methodologies, solidifying its status as a transformative paradigm in adaptive learning.
@@
Continual Learning is a critical field in artificial intelligence aimed at developing models that can learn new tasks while preserving knowledge from previous tasks. This field encompasses both task-specific strategies and generalization-based approaches. Task-specific strategies utilize four primary methodologies to address continual learning challenges: replay-based, parameter regularization, dynamic networks, and knowledge distillation. Replay-based methods [7, 80, 18, 22] mitigate catastrophic forgetting through selective sample storage and data generation. Regularization-based approaches [52, 4, 68, 97] limit alterations to crucial parameters, integrating various regularization methods to stabilize performance across tasks. Architecture-based strategies [69, 90, 89, 99] modify network structures to accommodate new information, employing dynamically expanding ensembles and structural adjustments based on task relevance. Distillation-based techniques [44, 46, 45, 33, 88] leverage knowledge transfer to maintain continuity of information across tasks. Generalization-based methods focus on model generalization that inherently support knowledge transfer and retention across tasks. [49] delves into balancing retention with generalization, while [64] provides insights into the interplay between learning new information and retaining old knowledge. [65] examines the role of controlled forgetting in enhancing model resilience, and [1] assesses the implications of network reinitialization on learning dynamics and generalization. Similar foundational investigations by [84, 37, 3, 16] explore continual learning through the lenses of NTK and generalization theory, though these studies primarily address traditional continual learning scenarios and do not fully integrate advancements from the era of pre-trained models.
@@
Parameter-Efficient Fine-Tuning for Continual Learning has established itself as an effective strategy to to counter catastrophic forgetting by training minimal additional parameters atop pre-trained models. Notable approaches such as L2P [82] and DualPrompt [81] introduce task-specific and dual prompts, respectively, facilitating adaptive task-specific learning while preserving invariant knowledge. S-Prompt [79] employs structural prompts to map discriminative domain relationships, while CODA-Prompt [71] applies Schmidt orthogonalization to refine these prompts. Concurrently, DAP [36] innovates by constructing real-time instance-level dynamic sub-networks to adapt to various domains. Moreover, HiDe-Prompt [77] integrates task-level knowledge sub-networks with distributional statistics to sample past data, thereby preventing suboptimal learning trajectories. EASE [100] further contributes by optimizing task-specific and expandable adapters to enhance knowledge retention. Despite these advancements, the reliance on specific configurations underscores the necessity for deeper theoretical exploration to fundamentally address PEFT-CL’s challenges, necessitating a shift towards a NTK perspective to enrich our understanding and efficacy in PEFT-CL.
@@ 3 Preliminaries
In the PEFT-CL context, we augment pre-trained models with adaptive sub-networks to manage sequential tasks. Let $f^{*}_{0}$ and $f^{*}_{T}$ denote the initial and target parameter spaces respectively, with * indicating optimized parameters. Given a series of tasks $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$, where each $\mathcal{D}_{\tau}$ comprises samples $(x,y)$ from $(X_{\tau},Y_{\tau})$, we introduce task-specific optimizable sub-network parameters $p_{\tau}$. The transformed model is represented as $f^{*}_{\tau}=(f^{*}_{0}\circ p_{\tau}\circ X_{\tau}\circ Y_{\tau})$, with $\circ$ denoting component integration. This configuration, inspired by L2P [82], features distinct class boundaries without explicit task identification during training, aligning with practical scenarios.
@@
$\frac{\partial f(x,\theta(t))}{\partial t}=-\Phi(x,X)\nabla_{f}\mathcal{L}(f(X,\theta(t)),Y)$ (2)
@@
This formulation elucidates the network’s trajectory towards the global minimum, exhibiting exponential convergence under a positive definite NTK [32, 42, 56, 8, 5, 92]. Furthermore, in PEFT-CL, to better adapt it for sequence learning scenarios, we have transformed it in Appendix A as follows:
@@
where $\Phi_{i}$ represents the locally converged NTK matrix for the $i$-th task.
@@ 4 Theoretical Insights
Harnessing the interpretative power of the NTK to decode network training dynamics, we assess the model’s resilience against forgetting through the generalization gaps and population losses. Initially, we derive the general formulation of cross-task generalization gap and population loss for the PEFT-CL scenario, addressing data from the $\tau$-th task post the final training session. We further extend our analysis, which assesses the population loss for individual tasks using NTK spectral theory. By examining the commonalities in these losses, we identify key elements that influence the optimization process of the PEFT-CL model and propose further theoretical insights. These concepts will be elaborated upon in a step-by-step manner.  [^1] The derivation process is thoroughly detailed in Appendix A, Appendix B, and Appendix C. We extend our appreciation to the contributions from [16, 10, 8, 3] for their invaluable assistance in theoretical derivations, some of which we reference in our work..
@@
To clarify the exposition, we detail the derivation processes for theorem 1 and theorem 2 in Appendix B and Appendix C, respectively. Building on these foundations, we further analyze and derive lemma 3, establishing the basis for the details of subsequent NTK-CL implementations.
@@
From lemma 3, we identify the key factors that need attention during the optimization process of the PEFT-CL model and further propose the NTK-CL framework, planned from the NTK perspective, based on these focal points.

### v2

@@ Abstract
Parameter-efficient fine-tuning for continual learning (PEFT-CL) has shown promise in adapting pre-trained models to sequential tasks while mitigating catastrophic forgetting problem. However, understanding the mechanisms that dictate continual performance in this paradigm remains elusive. To unravel this mystery, we undertake a rigorous analysis of PEFT-CL dynamics to derive relevant metrics for continual scenarios using Neural Tangent Kernel (NTK) theory. With the aid of NTK as a mathematical analysis tool, we recast the challenge of test-time forgetting into the quantifiable generalization gaps during training, identifying three key factors that influence these gaps and the performance of PEFT-CL: training sample size, task-level feature orthogonality, and regularization. To address these challenges, we introduce NTK-CL, a novel framework that eliminates task-specific parameter storage while adaptively generating task-relevant features. Aligning with theoretical guidance, NTK-CL triples the feature representation of each sample, theoretically and empirically reducing the magnitude of both task-interplay and task-specific generalization gaps. Grounded in NTK analysis, our framework imposes an adaptive exponential moving average mechanism and constraints on task-level feature orthogonality, maintaining intra-task NTK forms while attenuating inter-task NTK forms. Ultimately, by fine-tuning optimizable parameters with appropriate regularization, NTK-CL achieves state-of-the-art performance on established PEFT-CL benchmarks. This work provides a theoretical foundation for understanding and improving PEFT-CL models, offering insights into the interplay between feature representation, task orthogonality, and generalization, contributing to the development of more efficient continual learning systems.
@@ 1 Introduction
In practical applications, the relentless evolution of environments underscores the urgency for learning systems that can progressively accumulate knowledge. This has led to the prominence of Continual Learning (CL) [93, 18, 64, 47, 83, 59], a cornerstone task that equips the learning models with the ability to seamlessly assimilate fresh information over time, while mitigating catastrophic forgetting, i.e., a phenomenon that erodes previously acquired knowledge. In recent years, with the proliferation of pre-trained models possessing strong generalization capabilities [7, 68], researchers have discovered that they can empower early exploratory methods [8, 86, 23, 57, 5, 74, 102, 75, 96, 95, 107, 48, 49, 94], enabling CL systems to integrate new knowledge more efficiently. However, full fine-tuning of pre-trained models is computationally intensive and may compromise their original generalization capabilities [32, 99]. Thus, as a promising paradigm, Parameter-Efficient Fine-Tuning for Continual Learning (PEFT-CL) emerges as an alternative, updating only a minimal set of additional parameters while keeping the pre-trained model intact. Specifically, PEFT-CL not only offers a more philosophically sound framework akin to Socratic dialogue but also provides a lightweight training process that avoids generalization deterioration associated with full-scale fine-tuning [38, 81]. In addition, this seamless integration of new and old knowledge aligns with the wisdom expressed by Bernard of Chartres, demonstrating how PEFT-CL builds upon pre-existing knowledge to achieve a more adaptive learner with robust memory capabilities.
@@
Despite initial successes in mitigating catastrophic forgetting [88, 87, 76, 25, 109], PEFT-CL largely relies on subjective human insights and experiential doctrines for network design and enhancement, lacking a rigorous mathematical foundation. This reliance on non-theoretical intuition constrains the potential for a deeper understanding and advancement of the fundamental mechanisms within these learning systems. While Hide-Prompt [82] acknowledges the importance of addressing this issue and offers a loss-based perspective, it falls short of modeling optimization dynamics and pinpointing key factors. Therefore, to address this gap, we adopt the Neural Tangent Kernel (NTK) theory [36, 9, 6] as a robust mathematical tool to delve deeply into the intricacies of PEFT-CL optimization. Through this rigorous analysis, we derive several fundamental theorems and lemmas, including theorem 1, theorem 2, lemma 3, and theorem 4. While initially considered from a CL perspective, these have been generalized to the PEFT-CL scenario, providing profound insights into the key factors essential for effectively combating catastrophic forgetting in PEFT-CL optimization. Guided by these theories and key factors, we develop an NTK-CL framework, effectively reducing the quantified catastrophic forgetting discussed later.
@@
In addition to theoretical advantages, we also detail the differences in structure and optimization between our NTK-CL framework and current mainstream methodologies in Fig. 1. Unlike the Additional Subnetworks paradigm (Fig. 1a), which constructs task-specific subnetwork parameter spaces and concatenates features from all network parameter spaces at inference time [109, 25, 51], or the Prompts Optimization paradigm (Fig. 1b), which builds task-specific prompt pools for input interaction and employs cosine similarity for prompt selection [88, 39, 87, 76, 66], our NTK-CL (Fig. 1c) framework eliminates the need for task-specific parameter storage or prompt pools. Instead, it leverages a shared network parameter space across all tasks to adaptively generate task-relevant features based on input characteristics. Specifically, its design and optimization are entirely derived from NTK-based generalization gaps, which not only triple the sample representations but also consider knowledge retention, task-feature dissimilarity, and regularization term.
@@
(1) Theoretical Exploration of PEFT-CL: We pioneer the analysis of PEFT-CL through NTK lens and foundational mathematics. Through a series of derived theorems and lemmas, we identify critical factors that optimize PEFT-CL learners, including the number of samples in data subsets, the total sample volume across the dataset, knowledge retention strategies, task-feature dissimilarity constraints, and adjustments to regularization terms.
(2) Innovative Solutions Based on Key Factors: Closely aligned with the key factors derived from our theoretical analysis, we propose an NTK-CL framework specifically designed for the PEFT-CL scenario. First, guided by theorem 1 and theorem 2, to increase the sample size available for optimizing the PEFT-CL model without incurring excessive training costs, we incorporate multiple interventions to expand the representational breadth, ensuring that each sample is mapped to different spaces, effectively tripling the representational scope. Second, unlike most previous PEFT-CL methods that do not consider knowledge retention, we design an adaptive Exponential Moving Average (EMA) mechanism that preserves intra-task NTK forms in theorem 1, thereby enhancing knowledge retention. Additionally, we no longer focus on class-level orthogonality as in previous studies, but instead introduce task-feature orthogonality constraints that attenuate inter-task NTK forms in theorem 1, increasing knowledge separability. This dual approach not only effectively avoids the storage overhead associated with parameter preservation but also achieves superior continual performance. Finally, to ensure that network training aligns with the process of finding the saddle point solution in Eq. 32, we implement tailored regularization adjustments. These strategies optimally minimize the generalization gaps and population losses in both task-interplay and task-specific settings within the PEFT-CL scenario, mitigating the catastrophic forgetting problem both theoretically and practically.
(3) Empirical Validation on Diverse Datasets: We conduct extensive experiments across various datasets to validate the effectiveness of our key factors and methodologies. Additionally, we perform fair comparisons against numerous state-of-the-art methods, ensuring consistent task segmentations to mitigate performance discrepancies. This comprehensive validation substantiates the efficacy of our theoretical innovations in practical applications.
@@ 2 Related Works
Parameter-Efficient Fine-Tuning has emerged as a pivotal paradigm for optimizing model performance while mitigating computational and memory burdens associated with large-scale model adaptation. Seminal works introduce diverse methodologies, including Adapter modules [31], Low-Rank Adaptation (LoRA) [32], Prefix Tuning [50], Prompt Tuning [7], and BitFit [101]. These approaches demonstrate the efficacy of selectively fine-tuning components or introducing compact, trainable subnetworks within pre-trained architectures. Subsequent advancements further expand PEFT’s scope and capabilities. Jia et al. [37] pioneer efficient prompt tuning techniques for vision transformers, extending PEFT’s applicability to the visual domain. Zhou et al. [111] introduce contextual prompt fine-tuning, enhancing model adaptability while preserving generalization. Recent comprehensive studies [91, 92] reinforce PEFT’s critical role in enhancing model generalization and efficiency. These investigations rigorously analyze the theoretical underpinnings and empirical efficacy of various PEFT methodologies, solidifying its status as a transformative paradigm in adaptive learning.
@@
Continual Learning is a critical field in artificial intelligence aimed at developing models that can learn new tasks while preserving knowledge from previous tasks. In general, this field can be categorized into task-specific and generalization-based approaches. Task-specific strategies include four main methodologies: replay, regularization, dynamic architectures, and knowledge distillation. Replay methods [8, 86, 23] combat catastrophic forgetting by storing or generating representative samples. Regularization techniques [57, 5, 74, 102] constrain changes to critical parameters, ensuring stability across tasks. Dynamic architectures [75, 96, 95, 107] adapt network structures to incorporate new information, often through expansion or task-relevant modifications. Knowledge distillation [48, 49, 94] transfers learned knowledge, maintaining information continuity. Generalization-based methods emphasize intrinsic model capabilities for knowledge transfer and retention. Lin et al. [52] investigate the balance between retention and generalization. Raghavan et al. [69] analyze the interaction between learning new information and preserving old knowledge. Ramkumar et al. [70] study controlled forgetting to enhance model robustness, while Alabdulmohsin et al. [1] examine the effects of network reinitialization on learning and generalization. Additional foundational research [90, 40, 4, 20] explores CL through the lenses of NTK and generalization theory, though these studies primarily address traditional continual learning scenarios and do not fully integrate advancements from the era of pre-trained models.
@@
Parameter-Efficient Fine-Tuning for Continual Learning has established itself as an effective strategy to counter catastrophic forgetting by training minimal additional parameters atop pre-trained models. Notable approaches such as L2P [88] and DualPrompt [87] introduce task-specific and dual prompts, respectively, facilitating adaptive task-specific learning while preserving invariant knowledge. S-Prompt [85] employs structural prompts to map discriminative domain relationships, while CODA-Prompt [76] applies Schmidt orthogonalization to refine these prompts. In parallel, DAP [39] proposes the construction of real-time, instance-level dynamic subnetworks, offering a flexible mechanism to accommodate the nuances of diverse domains. HiDe-Prompt [82] integrates hierarchical task-level knowledge subnetworks with distributional statistics to sample past data, effectively curbing suboptimal learning trajectories. EASE [109] further contributes by optimizing task-specific, expandable adapters, thereby fortifying the model’s capacity for knowledge retention. Despite these significant strides, the reliance on particular configurations highlights the imperative for a more profound theoretical investigation to fundamentally tackle the challenges inherent in PEFT-CL. This necessitates a paradigm shift toward a NTK perspective, which promises to enrich our understanding in PEFT-CL.
@@ 3 Preliminaries
In the PEFT-CL context, we augment pre-trained models with adaptive subnetworks to manage sequential tasks. Let $f^{*}_{0}$ and $f^{*}_{T}$ denote the initial and target parameter spaces respectively, with * indicating optimized parameters. Given a series of tasks $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$, where each $\mathcal{D}_{\tau}$ comprises samples $(x,y)$ from $(X_{\tau},Y_{\tau})$, we introduce task-specific optimizable subnetwork parameters $p_{\tau}$. The transformed model is represented as $f^{*}_{\tau}=(f^{*}_{0}\circ p_{\tau}\circ X_{\tau}\circ Y_{\tau})$, with $\circ$ denoting component integration. This configuration, inspired by L2P [88], features distinct class boundaries without explicit task identification during training, aligning with practical scenarios.
@@
$\frac{\partial f(x,\theta(t))}{\partial t}=-\Phi(x,X)\nabla_{f}\mathcal{L}(f(X,\theta(t)),Y),$ (2)
@@
This formulation elucidates the network’s trajectory towards the global minimum, exhibiting exponential convergence under a positive definite NTK [36, 46, 9, 6, 97]. Furthermore, in PEFT-CL, to better adapt it for sequence learning scenarios, we have transformed it in Appendix B as follows:
@@
where $\Phi_{i}$ denotes the locally converged NTK matrix for the $i$-th task, and $\lambda$ is the hyper-parameter that controls the L2 regularization of the trainable parameters in Eq. 32. This hyper-parameter is crucial for finding the dynamic saddle point solution of the model in PEFT-CL scenario.
@@ 4 Theoretical Insights
Harnessing the interpretative power of the NTK to decode network training dynamics, we assess the model’s resilience against forgetting through the generalization gaps and population losses. Initially, we derive the general formulation of cross-task generalization gap and population loss for the PEFT-CL scenario, addressing data from the $\tau$-th task post the final training session. We further extend our analysis, which assesses the population loss for individual tasks using NTK spectral theory. By examining the commonalities in these losses, we identify key elements that influence the optimization process of the PEFT-CL model and propose further theoretical insights. These concepts will be elaborated upon in a step-by-step manner.  [^1] The derivation process is thoroughly detailed in Appendix B, Appendix C, and Appendix D. We extend our appreciation to the contributions from [20, 11, 9, 4] for their invaluable assistance in theoretical derivations, some of which we reference in our work..
@@
To clarify the exposition, we detail the derivation processes for theorem 1 and theorem 2 in Appendix C and Appendix D, respectively. Building on these foundations, we further analyze and derive lemma 3, establishing the basis for the details of subsequent NTK-CL implementations.
@@
From lemma 3, we identify the key factors that require attention during the optimization process of the PEFT-CL model and propose the NTK-CL framework. While these key factors may also play a beneficial role in other paradigms or traditional CL approaches, our NTK-CL framework introduces specific improvements and innovations tailored for the PEFT-CL scenario. Each component is meticulously designed to align with the constraints and requirements derived from our theoretical analysis, thereby addressing the limitations of existing PEFT-CL methods.